# Hadamard-type formulas for real eigenvalues of canonically symplectic operators

### v2

@@ 4.1. Maslov crossing forms
Towards the proof of Theorem 4.2, we begin by computing crossing forms. The following discussion mimics that in [LS20first, §4.5]. Recall from (3.3) that
@@
The result follows from the fact that the operator $(N^{*}+V_{t}^{*}-\lambda)\tau=(\tau N)^{*}+V_{t}^{*}\tau-\lambda\tau$ is a bounded perturbation of the adjoint $(\tau N)^{*}$ of a symmetric operator $\tau N$ satisfying the Green’s identity (2.21) with the trace operator $\operatorname{T}=\operatorname{T}_{t}$ and [BBvnkF98, Proposition 3.5]. ∎
@@
see, e.g. [CJLS14, Lemma 3.8]. Then $(\lambda_{0},t_{0})$ is called a conjugate point if $\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}\neq\{0\}$ (or either of the equivalent assertions in (4.5) hold). The $t$-crossing form for the Lagrangian plane $t\mapsto\Upsilon_{\lambda_{0},t}$ with respect to $\mathfrak{D}$ on the finite dimensional intersection $\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}$ is defined to be
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{p})\coloneqq\frac{d}{dt}\big|_{t=t_{0}}\tilde{\Omega}(\mathbf{q},\mathscr{M}_{\lambda_{0},t}\mathbf{p}),\qquad\mathbf{q},\mathbf{p}\in\Upsilon_{\lambda_{0},t_{0}},$ (4.7)
@@
Finally, following [PT09], when $\mathfrak{m}_{\lambda_{0}}=0$ we define the second order crossing form to be
$\mathfrak{m}^{(2)}_{\lambda_{0}}(\mathbf{q},\mathbf{p})=\frac{d^{2}}{d\lambda^{2}}\big|_{\lambda=\lambda_{0}}\tilde{\Omega}(\mathbf{q},\mathscr{M}_{\lambda,t_{0}}\mathbf{p}).$ (4.9)
The following is an analogue of [LS20first, Lemma 4.19].
@@
We begin with the proof of (4.10) and the statements preceding those equations. The proof is almost identical to that of [LS20first, Lemma 4.19], but we give it here for completeness. Let us momentarily denote $K_{t}\coloneqq\ker\left((N^{*}+V_{t}-\lambda)\tau\right)$, so that $\mathbb{K}_{\lambda,t}=\operatorname{T}_{t}K_{t}$. We also denote $q\coloneqq\operatorname{T}_{t_{0}}\mathbf{u}$, $\mathbf{q}\coloneqq(q,q)$, and let $\mathbb{P}_{t}$ be the orthogonal projections onto $\mathbb{K}_{\lambda,t}$. Then $\mathbb{P}_{t}\in C^{1}(\mathcal{I},\mathcal{B}(\mathfrak{H}^{4}))$ for some open interval $\mathscr{I}\subseteq\mathcal{I}$ centred at $t_{0}$ (see, for example, [BBvnkF98, Theorem 3.9]).
@@
It remains to show that the mapping $t\mapsto\mathbf{w}_{t}$ is in $C^{1}(\mathscr{I},\mathcal{H}_{+}^{2})$; we will exploit the bijectivity of the mapping (4.15), although some care is needed to handle the $t$-dependent domain of that mapping. Let $U_{t}$ denote the $C^{1}$ family of boundedly invertible transformation operators in $\mathcal{H}_{+}^{2}$ that split the projections $\mathcal{P}_{K_{t}}$ onto $K_{t}$ and $\mathcal{P}_{K_{t_{0}}}$ onto $K_{t_{0}}$ such that $U_{t}\mathcal{P}_{K_{t_{0}}}=\mathcal{P}_{K_{t}}U_{t}$ and $U_{t}:K_{t_{0}}\to K_{t}$ are bijections for $t$ near $t_{0}$, see [LSS18, Remark 2.4], [CJLS14, Remark 3.5], [Dalecki_Krein, Section IV.1], [F04, Remark 6.11]. Introducing $\mathbf{v}_{t}\in K_{t_{0}}$ by $\mathbf{v}_{t}=U_{t}^{-1}\mathbf{w}_{t}$ so that $\operatorname{T}_{t}\mathbf{w}_{t}=h_{t}$ yields $(\operatorname{T}_{t}\circ U_{t})\mathbf{v}_{t}=h_{t}$. The map $\operatorname{T}_{t}\circ U_{t}|_{K_{t_{0}}}:K_{t_{0}}\to\operatorname{ran}\mathbb{P}_{t}$ is a bijection and $t\mapsto\operatorname{T}_{t}\circ U_{t}|_{K_{t_{0}}}$ is in $C^{1}(\mathcal{I},\mathcal{B}(K_{t_{0}},\mathfrak{H}^{4}))$ by the assumptions in the lemma. Since $\mathbf{w}_{t}=U_{t}\circ(\operatorname{T}_{t}\circ U_{t})^{-1}h_{t}$, the function $t\mapsto\mathbf{w}_{t}$ is $C^{1}$ because each of the terms in the composition is $C^{1}$.
@@
4.2. Lyaponuv-Schmidt reduction
@@
Our goal is to construct a matrix $M(\lambda,t)$, the zero set of the determinant of which locally coincides with the real spectrum of $\mathcal{N}_{t}+V_{t}$ (i.e. for $(\lambda,t)$ near $(\lambda_{0},t_{0})$). We proceed with Lyaponuv-Schmidt reduction.
@@
$\mathbf{u}=\sum_{i=1}^{n}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}+\tilde{\mathbf{u}}$, (4.34)
@@ 5. Application: Linearised NLS for standing waves on compact star graphs
The minimal symmetric operator with finite and equal deficiency indices acting in $\mathcal{H}^{2}=\left(L^{2}(\mathcal{G})\right)^{2}$ is then $A=-\partial_{xx},\,\operatorname{dom}(A)=\widehat{H}_{0}^{2}(\mathcal{G})$, with maximal adjoint operator $A^{*}=-\partial_{xx},\,\operatorname{dom}(A^{*})=\mathcal{H}_{+}=\widehat{H}^{2}(\mathcal{G})$. The trace operators
@@
Indeed such a choice is possible, since for each $t\in[\varepsilon_{0},1]$, $\mathcal{N}_{t}+V_{t}$ is a bounded perturbation of a skew self-adjoint operator, and therefore its spectrum is contained in the vertical strip $\left\{z\in\mathbb{C}:|{\rm{Re}}(z)|\leq\|V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}\right\}$ around the imaginary axis in the complex plane [Kato]. Taking the supremum of the spectral bounds over $t\in[\varepsilon_{0},1]$ yields the claim.
If $\varepsilon_{0}>0$ is small enough, we claim that $\operatorname{Spec}(\mathcal{N}+V_{\varepsilon_{0}})\cap\mathbb{R}=\emptyset$; the proof is similar to that of [CCLM23, Lemma 3.23]. First, we note that the operators $\mathcal{A}+F_{\varepsilon_{0}}$ and $\mathcal{A}+G_{\varepsilon_{0}}$ with domain $\operatorname{dom}(\mathcal{A})=\{u\in\mathcal{H}_{+}:\operatorname{tr}_{t}u\in\mathcal{L}\}$ are strictly positive. To see this, since $\alpha>0$ we note that $\operatorname{Spec}(\mathcal{A})\subset(0,\infty)$, hence, $\mathcal{A}\geq C$ for sufficiently small $C>0$. Using this inequality we arrive at
@@
for sufficiently small $c>0$, where $g(x)=(2p+1)\phi(x)+\beta$. A similar inequality holds for $\mathcal{A}+F_{\varepsilon_{0}}$. Next, suppose, by way of contradiction, that there exists $\varepsilon_{0}^{2}\lambda\in\operatorname{Spec}(\mathcal{N}_{\varepsilon_{0}})\cap\mathbb{R}$ with eigenfunction $\mathbf{u}_{t_{0}}=(u_{t_{0}},v_{t_{0}})\in\operatorname{dom}(\mathcal{N}_{\varepsilon_{0}})$. Using the fact that $\mathcal{A}+G_{\varepsilon_{0}}$ is strictly positive and arguing as in [CCLM23, Lemma 3.21] we note that the eigenvalue problem (1.10) for $t=\varepsilon_{0}$ small enough is equivalent to the self-adjoint eigenvalue problem

### v3

@@ 4.1. Maslov crossing forms
Towards the proof of Theorem 4.2, we begin by computing crossing forms. The following discussion mimics that in [LS24, §4.5]. Recall from (3.3) that
@@
The result follows from the fact that the operator $(N^{*}+V_{t}^{*}-\lambda)\tau=(\tau N)^{*}+V_{t}^{*}\tau-\lambda\tau$ is a bounded perturbation of the adjoint $(\tau N)^{*}$ of a symmetric operator $\tau N$ satisfying the Green’s identity (2.21) with the trace operator $\operatorname{T}=\operatorname{T}_{t}$ and [BF98, Proposition 3.5]. ∎
@@
see, e.g. [CJL+16, Lemma 3.8]. Then $(\lambda_{0},t_{0})$ is called a conjugate point if $\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}\neq\{0\}$ (or either of the equivalent assertions in (4.5) hold). The $t$-crossing form for the Lagrangian plane $t\mapsto\Upsilon_{\lambda_{0},t}$ with respect to $\mathfrak{D}$ on the finite dimensional intersection $\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D}$ is defined to be
$\mathfrak{m}_{t_{0}}(\mathbf{q},\mathbf{p})\coloneqq\frac{d}{dt}\big|_{t=t_{0}}\tilde{\Omega}(\mathbf{q},\mathscr{M}_{\lambda_{0},t}\mathbf{p}),\qquad\mathbf{q},\mathbf{p}\in\Upsilon_{\lambda_{0},t_{0}}\cap\mathfrak{D},$ (4.7)
@@
The following is an analogue of [LS24, Lemma 4.19].
@@
We begin with the proof of (4.9) and the statements preceding those equations. The proof is almost identical to that of [LS24, Lemma 4.19], but we give it here for completeness. Let us momentarily denote $K_{t}\coloneqq\ker\left((N^{*}+V_{t}-\lambda)\tau\right)$, so that $\mathbb{K}_{\lambda,t}=\operatorname{T}_{t}K_{t}$. We also denote $q\coloneqq\operatorname{T}_{t_{0}}\mathbf{u}$, $\mathbf{q}\coloneqq(q,q)$, and let $\mathbb{P}_{t}$ be the orthogonal projections onto $\mathbb{K}_{\lambda,t}$. Then $\mathbb{P}_{t}\in C^{1}(\mathcal{I},\mathcal{B}(\mathfrak{H}^{4}))$ for some open interval $\mathscr{I}\subseteq\mathcal{I}$ centred at $t_{0}$ (see, for example, [BF98, Theorem 3.9]).
@@
It remains to show that the mapping $t\mapsto\mathbf{w}_{t}$ is in $C^{1}(\mathscr{I},\mathcal{H}_{+}^{2})$; we will exploit the bijectivity of the mapping (4.14), although some care is needed to handle the $t$-dependent domain of that mapping. Let $U_{t}$ denote the $C^{1}$ family of boundedly invertible transformation operators in $\mathcal{H}_{+}^{2}$ that split the projections $\mathcal{P}_{K_{t}}$ onto $K_{t}$ and $\mathcal{P}_{K_{t_{0}}}$ onto $K_{t_{0}}$ such that $U_{t}\mathcal{P}_{K_{t_{0}}}=\mathcal{P}_{K_{t}}U_{t}$ and $U_{t}:K_{t_{0}}\to K_{t}$ are bijections for $t$ near $t_{0}$, see [LSS18, Remark 2.4], [CJL+16, Remark 3.5], [DK74, Section IV.1], [FUR04, Remark 6.11]. Introducing $\mathbf{v}_{t}\in K_{t_{0}}$ by $\mathbf{v}_{t}=U_{t}^{-1}\mathbf{w}_{t}$ so that $\operatorname{T}_{t}\mathbf{w}_{t}=h_{t}$ yields $(\operatorname{T}_{t}\circ U_{t})\mathbf{v}_{t}=h_{t}$. The map $\operatorname{T}_{t}\circ U_{t}|_{K_{t_{0}}}:K_{t_{0}}\to\operatorname{ran}\mathbb{P}_{t}$ is a bijection and $t\mapsto\operatorname{T}_{t}\circ U_{t}|_{K_{t_{0}}}$ is in $C^{1}(\mathcal{I},\mathcal{B}(K_{t_{0}},\mathfrak{H}^{4}))$ by the assumptions in the lemma. Since $\mathbf{w}_{t}=U_{t}\circ(\operatorname{T}_{t}\circ U_{t})^{-1}h_{t}$, the function $t\mapsto\mathbf{w}_{t}$ is $C^{1}$ because each of the terms in the composition is $C^{1}$.
@@
4.2. Lyapunov-Schmidt reduction
@@
Our goal is to construct a matrix $M(\lambda,t)$, the zero set of the determinant of which locally coincides with the real spectrum of $\mathcal{N}_{t}+V_{t}$ (i.e. for $(\lambda,t)$ near $(\lambda_{0},t_{0})$). We proceed with Lyapunov-Schmidt reduction.
@@
$\mathbf{u}=\sum_{i=1}^{g}\alpha_{i}\mathbf{u}_{t_{0}}^{(i)}+\tilde{\mathbf{u}}$, (4.33)
@@ 5. Application: Linearised NLS for standing waves on compact star graphs
The minimal symmetric operator with finite and equal deficiency indices acting in $\mathcal{H}^{2}=\left(L^{2}(\mathcal{G})\right)^{2}$ is then $A=-\partial_{xx},\,\operatorname{dom}(A)=\widehat{H}_{0}^{2}(\mathcal{G})$, with maximal adjoint operator $A^{*}=-\partial_{xx},\,\operatorname{dom}(A^{*})=\mathcal{H}_{+}=\widehat{H}^{2}(\mathcal{G})$.
The trace operators
@@
Indeed such a choice is possible, since for each $t\in[\varepsilon_{0},1]$, $\mathcal{N}_{t}+V_{t}$ is a bounded perturbation of a skew self-adjoint operator, and therefore its spectrum is contained in the vertical strip $\left\{z\in\mathbb{C}:|{\rm{Re}}(z)|\leq\|V_{t}\|_{\mathcal{B}(\mathcal{H}^{2})}\right\}$ around the imaginary axis in the complex plane [KAT80]. Taking the supremum of the spectral bounds over $t\in[\varepsilon_{0},1]$ yields the claim.
If $\varepsilon_{0}>0$ is small enough, we claim that $\operatorname{Spec}(\mathcal{N}+V_{\varepsilon_{0}})\cap\mathbb{R}=\emptyset$; the proof is similar to that of [CCL+23, Lemma 3.23]. First, we note that the operators $\mathcal{A}+F_{\varepsilon_{0}}$ and $\mathcal{A}+G_{\varepsilon_{0}}$ with domain $\operatorname{dom}(\mathcal{A})=\{u\in\mathcal{H}_{+}:\operatorname{tr}_{t}u\in\mathcal{L}\}$ are strictly positive. To see this, since $\alpha>0$ we note that $\operatorname{Spec}(\mathcal{A})\subset(0,\infty)$, hence, $\mathcal{A}\geq C$ for sufficiently small $C>0$. Using this inequality we arrive at
@@
for sufficiently small $c>0$, where $g(x)=(2p+1)\phi(x)+\beta$. A similar inequality holds for $\mathcal{A}+F_{\varepsilon_{0}}$. Next, suppose, by way of contradiction, that there exists $\varepsilon_{0}^{2}\lambda\in\operatorname{Spec}(\mathcal{N}_{\varepsilon_{0}})\cap\mathbb{R}$ with eigenfunction $\mathbf{u}_{t_{0}}=(u_{t_{0}},v_{t_{0}})\in\operatorname{dom}(\mathcal{N}_{\varepsilon_{0}})$. Using the fact that $\mathcal{A}+G_{\varepsilon_{0}}$ is strictly positive and arguing as in [CCL+23, Lemma 3.21] we note that the eigenvalue problem (1.10) for $t=\varepsilon_{0}$ small enough is equivalent to the self-adjoint eigenvalue problem